OpenStudy (anonymous):

10(x – 3) = 5(3x – 8) @StudyGurl14

OpenStudy (studygurl14):

Okay, so this one might look a little trickier, but it really isn't. All you have to do is apply the distributive property.

OpenStudy (anonymous):

Okay i tried but i don't come up with the same answers as the 4 listed

OpenStudy (anonymous):

A. x = 2 B. x = 10 C. x = –2 D. x = –10

OpenStudy (anonymous):

c: -2 d: -10

OpenStudy (studygurl14):

Okay, so your first step would be to distribute the 10 to the x and -3 inside the parentheses. You would get what?

OpenStudy (anonymous):

10x and -30

OpenStudy (anonymous):

I thought it was 10x - 30 to begin with how do you tell the difference?

OpenStudy (studygurl14):

correct

OpenStudy (studygurl14):

both are correct. So now you have, 10x - 30 = 5(3x - 8)

OpenStudy (studygurl14):

You can distribute the 5 to the 3x -8 as well. What would that give you?

OpenStudy (anonymous):

15x and -40

OpenStudy (studygurl14):

right

OpenStudy (studygurl14):

so now you have 10x - 30 = 15x - 40 now we have a very similar problem to what we did last time. What's the next step?

OpenStudy (anonymous):

isolateing the variable

OpenStudy (studygurl14):

true. we do need to isolate the variable. What's the next step towards achieving that?

OpenStudy (anonymous):

25x - 30 = -40 right?

OpenStudy (anonymous):

This is the part that gets me when i add the variables.

OpenStudy (anonymous):

I don't know which to add too.

OpenStudy (studygurl14):

Um...not exactly.

OpenStudy (studygurl14):

Here's what we'll do. We'll move the -30 to the right side, and the 15x to the left side, ok?

OpenStudy (studygurl14):

What's the opposite of -30?

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

30

OpenStudy (studygurl14):

right

OpenStudy (studygurl14):

so you want to add 30 to both sides of the equation. What do you get?

OpenStudy (anonymous):

10x - 70 = -15x

OpenStudy (studygurl14):

Um...no.

OpenStudy (studygurl14):

10x - 30 = 15x - 40 + 30 + 30 _____________________ 10x + 0 = 15x - 10

OpenStudy (studygurl14):

Which simplifies into 10x = 15x - 10

OpenStudy (studygurl14):

Now, we want to move the 15x to the left side of the equals sign. How do we do that?

OpenStudy (anonymous):

combining both like terms of 10 and 15x

OpenStudy (studygurl14):

right. First though, what's the opposite of 15x?

OpenStudy (anonymous):

-15x

OpenStudy (studygurl14):

correct. we need to subtract 15x from both sides of the equation in order to move the 15x from the right to the left. This gives you: 10x - 15x = 0 - 10 -5x = -10

OpenStudy (studygurl14):

What now?

OpenStudy (anonymous):

10 divided by -5

OpenStudy (studygurl14):

-10 divided by -5, correct. Which is?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

Thing is i was assigned this without being explained all these proccesses so i had to learn it all from a really old text book with some pages missing sorry for troubling you.

OpenStudy (studygurl14):

It's fine. You aren't troubling me. So then the answer is x = 2